Please show step by step to solve problem

Let's assume you sold short 200 shares of common stock at $60 per share. The initial margin is 50%. What would be the maintenance margin if a margin call was made at a stock price of $72?
Answer #1

shares of common stock=200
Price per share=$60

stock price = $72

The initial margin is 50%=0.50

Equity=  (shares of common stock*Price per share)*initial margin +((shares of common stock*Price per share)

=(200*60)*0.50 +(200*60)

=12,000*0.50 +12,000

=18,000

Maintenance Margin=(Equity - (shares of common stock*stock price))/ (shares of common stock*stock price)

=(18,000-200(72))/200(72)

=(18,000-14,400)/14,400

=0.25

or 25%

Maintenance Margin=25%